OpenStudy (anonymous):

Find the value of x. log1/2(1/4) = x

OpenStudy (anonymous):

x =2

OpenStudy (anonymous):

\[log_{\frac{1}{2}}(\frac{1}{4})=x\] means \[(\frac{1}{2})^x=\frac{1}{4}\] and since \[\sqrt{\frac{1}{4}}=(\frac{1}{4})^{\frac{1}{2}}\], \[x=\frac{1}{2}\]

OpenStudy (anonymous):

oops i meant since \[(\frac{1}{2})^2=\frac{1}{4}\] \[x=2\]